OpenStudy (anonymous):

integral 1 / [(x^2)sqrt(16-x^2)]. show steps/explain method.

OpenStudy (turingtest):

looks like a trig sub\[\sqrt{a^2-b^2x^2}\implies x=\frac ab\sin\theta\]

OpenStudy (anonymous):